I need some assistance figuring out this problem. Any feedback is appreciated.
Evaluate
((((( 24-[2(3-8)+2^3]+13+11(23-4^2/20-2x5 )))))

Answer by stanbon(48537)

24-[2(3-8)+2^3]+13+11(23-4^2/20-2x5
There is a closeed parenthesis missing toward the end of your problem.
=24-[-10+8]+13+11(23-(4/5)-10
=24-(-2)+13+11(13-(4/5))
=26+13+11(61/5)
=26+13+134+1/5
=173 1/5
Cheers,
Stan H
Comment: I don't think this is the answer you are looking for.
The missing parenthesis is important.